How to find the focus and directrix of 3x^2=-y
Answer by lwsshak3(11628)

How to find the focus and directrix of
3x^2=-y
x^2=-(1/3)y
This is an equation of a parabola that opens downward.
Its basic form: (x-h)^2=4p(y-k), (h,k)=(x,y) coordinates of the vertex
vertex:(0,0)
axis of symmetry: x=0 or y-axis
4p=1/3
p=1/12
focus: (0,-1/12)
directrix: y=1/12